how to find the equation of the circle having AB as diameter where A is the point (1,8) and b is the point (3,14)
Answer by checkley77(12844)

d^2=(3-1)^2+(14-8)^2
D^2=2^2+6^2
D^2=4+36
D^2=40
D=SQRT40
D=6.32 ANS.